OpenStudy (anonymous):

Very Good Morring to You students Just a quickquestion. Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together? A. 12 minutes B. 15 minutes C. 21 minutes D. 23 minutes E. 28 minutes

OpenStudy (anonymous):

Would you please explain the way that you get the result.

OpenStudy (anonymous):

The way to work this out is: \begin{eqnarray*} \frac{1}{T} &=& \frac{1}{30}+\frac{1}{45}+\frac{1}{90} \\ \frac{1}{T} &=& \frac{3+2+1}{90} \\ \frac{1}{T} &=& \frac{6}{90} \\ \frac{1}{T} &=& \frac{1}{15} \\ T &=& 15 \end{eqnarray*} Thereforer the answer is B. 15 minutes.